Chromium metal crytallises with a b.c.c lattice . The length of unit cell edge is 287pm . What is the atomic radius? 

Grade:

1 Answers

Sona Srivastava
13 years ago

imagine the main body diagonal of the cube -
there will b 3 atoms with their centres on the body diagonal

one atom(at centre) completely lies in the cube (contributing =2*r) and other 2 lie half inside the cube(contributing r each)

so

4*r = a√3 = length of body diagonal

4*r = 287√3 pm

r = (287√3)/4 pm

r = 124.27 pm